### v1

# Multi-messenger signatures of a deformed magnetar in gamma-ray bursts

Abstract
We study the evolution of a newly formed magnetized neutron-star (NS) as a power source of gamma-ray bursts (GRBs) in the light of both gravitational wave (GW) and electromagnetic (EM) radiations. The compressible and incompressible fluids are employed in order to model the secular evolution of Maclaurian spheroids. It is shown that the GW and EM light curves evolve as a function of eccentricity and rotational frequency with time. We find that the light curve characteristics crucially depend on NS parameters such as magnitude and structure of magnetic field, ellipticity and the equation of state (EOS) of the fluid. The presence of X-ray flares, whose origins are not yet well understood, can be captured in our model regarding some specific nuclear EOSs. Our model allowing us to explain flares that occur within the wide range of $10$ to $10^{4}$ seconds and the peak luminosity in the order of $10^{46}$ $10^{51}$ $\rm\text{erg}/s$ by using a reasonable set of parameters such as magnetic field strength around $10^{14}-10^{16}$ Gauss, the quadrupole to dipole ratio of magnetic field up to 500. By applying our model to a sample of GRB X-ray flares observed by Swift/XRT, we try to constraint the crucial parameters of a deformed magnetar via MCMC fitting method. Our analysis shows that ongoing and upcoming joint multi-messenger detections can be used to understand the nature of GRB’s central engine and its evolution at the early times of the burst formation.
Multi-messenger–Magnetar – Gamma-ray bursts – Gravitational Wave–High-Energy Emission
1 Introduction
The physics behind GRB’s central engine is still among astrophysics open questions, while many scenarios have been introduced to explain it (e.g., Li et al., 2018). The collapse of a massive compact star is a candidate for long duration GRB’s central engine Lü_2015 and compact binary mergers can be served as a source of short duration GRBs Song & Liu (2023). There are several evidences that the GRB and its afterglow may be powered by a rapidly rotating highly-magnetised NS in which lose of angular momentum is a result of EM and GW emissions or is due to the accretion processes Rowlinson et al. (2013); Mösta et al. (2020); Bucciantini (2012); Bianco et al. (2024). The plateau phase in X-ray afterglow emission is observed in many GRBs (e.g., Liang et al., 2007; Li et al., 2012, 2015; Dainotti et al., 2020, 2022, and references therein) and can be explained by the spin-down of an ultra-magnetized millisecond NS Stratta et al. (2018); Gompertz et al. (2013); Wang et al. (2024). It has been demonstrated that the spin-down luminosity from multipolar EM emission can effectively account for the plateau phase in the majority of GRBs afterglows offering the complex magnetic field structure of magnetars as one of the candidates for GRB’s central engine Wang et al. (2024).
X-ray flares are also frequent events occurring in about 33 percent of GRBs within a time interval ranging from 30 to $10^{5}$ seconds after GRB trigger time. The appearance of early X-ray flares during the light curves has been widely considered previously Ruffini et al. (2018b); Chincarini et al. (2010); Zheng et al. (2022); Saji et al. (2023); Zhang et al. (2006). It has been suggested that the prolonged activity of the central engine in GRB is responsible for flaring features alongside afterglow emissions Burrows et al. (2005); Fan & Wei (2005); Beniamini & Kumar (2016). However explaining the late time flaring activities is challenging in terms of black hole (BH) central engine and are more in favour of the idea of rapidly rotating magnetar central engines for GRBs Rowlinson et al. (2010); Metzger et al. (2008); Zhang & Mészáros (2001); Dai & Lu (1998). It has been investigated that the magneto hydrodynamic instabilities in a rotating neutron star can produce a strong temporary magnetic field of order $10^{17}G$, and the magnetic re-connection effects leads to multiple X-ray flares Kluźniak_1998; Dai et al. (2006). Beyond gamma-ray and X-ray observations, there are various data from optical to radio wavelengths which supports the magnetar central engine model Rowlinson et al. (2013); Xie et al. (2020); Lazarus et al. (2012). These observations offer high values of the magnetic field of the order of $B\sim 10^{12}G-10^{17}G$ with multipolar structures Wang et al. (2024). High values of quadrupole magnetic field component up to $\sim 10^{3}$ times greater than dipole magnetic field have been proposed to explain GRB features Ruffini et al. (2018a); Stratta et al. (2018).
The multi-messenger nature of magnetars through GW observations provides constraints on the deformation parameter, ellipticity, to be of order $\sim 10^{-5}-10^{-2}$asky & Glampedakis (2016). The GW signal generated by magnetars can be explored more precisely by current and upcoming GW detectors Lai & Shapiro (1995); Sur & Haskell (2021); Xie et al. (2022a). These signals are generated primarily due to the extreme magnetic fields and rotational instabilities that cause deviations from axisymmetry. Such observations will provide valuable insights into the physics of dense matter and the strong gravitational fields characteristic of these highly magnetized NSs. The polarization characteristics of the EM counterpart of GW created by coalescence of the binary sources in short GRBs has been considered previously in Shakeri & Allahyari (2018).
In this paper, we consider the evolution of a deforemed magnetar as a power source for GRBs. We use maclaurin spheroid to model a newborn magnetar to explain different properties of both EM and GW observations of GRBs. Maclaurin spheroid as a homogeneous rotating object was first introduced by S.Chandrasekhar (1969) and later on used to model the spin down of pulsars by Finn & Shapiro (1990), where the the evolution of eccentricity leads to spin variation of NS. Despite previous studies Xie et al. (2022a); Lasky & Glampedakis (2016); Suvorov & Kokkotas (2020); Rueda et al. (2022) in which the time evolution of angular velocity is derived directly from rotational energy loss of magnetar, here, we examine angular momentum as a function of the Maclaurin spheroid’s eccentricity, which evolves over time. Our approach is more fundamental since rotational evolution is originating from the deformation of the new NS which can result in both spin down and spin up of the object. It is demonstrated that in the absence of accretion, during the loss of angular momentum, new born NS can also experience spin-up Finn & Shapiro (1990), it is argued that this process eventually leads to the formation of a BH. However, in this paper, we examine the stability regions within the parameter space more accurately where the spin-up is caused by the decreasing moment of inertia due to the softness of the EoS, and BH formation does not occur.
We investigate the impact of key model parameters such as the magnetic field, quadrupole to dipole ratio, mass and EOS on both EM and GW light curves. We assume that both dipole and quadrupole emissions contribute to the EM radiation of the magnetar. We evaluate different features of the X-ray light curves under influence of the properties of the system such as the total mass of the progenitor, the magnetic field of the newborn magnetars for both compressible and incompressible fluids. Moreover, in the case of compressible fluids, we consider the impact of various polytropic (adiabatic) indices on the luminosity features. We see that for some specific values of the polytropic index a peak in the luminosity appears which will be faded out with time, this can provide an explanation for the X-ray flares in GRBs.
The present paper is organized as following. In Sec. 2, we review the main features and properties of the maclaurin spheroid in the context of both incompressible and compressible fluids. Then, we consider the evolutionary equation for the eccentricity due to the variation in angular momentum for both types of mentioned fluids. In Sec. 3, we discuss the evolution of NS’s dipole, quadrupole and total EM luminosity due to evolution of eccentricity caused by angular momentum loss. In Sec. 4, we evaluate the GW radiations of a new born NS and examine the detectability of resulting GW emission in the light of current and future detectors. In Sec. 5, we present total energy loss of the magnetar during spin down due to the evolution of the eccentricity and finally In Sec. 7, we discuss our results and give some concluding remarks. In this paper, we are working in geometrical unit in which $G=c=\hbar=1$.
2 FRAMEWORK
In this section, we present maclaurin spheroid as a model for the central engine of GRBs. The angular velocity of maclaurin spheroid is a function of eccentricity which measures the deformation of NS in polar plane and is defined as S.Chandrasekhar (1969); Finn & Shapiro (1990) :
$e=(1-\frac{c^{2}}{a^{2}})^{1/2},$ (1)
where a and c are shown in Fig.1 with c coinciding the rotational axis of neutron star. In order to consider GW emissions, we assume an triaxial configuration of NS with small deformation in equatorial plane so that ellipticity is defined as
$\epsilon$ $=\frac{a-b}{(a+b)/2},$ (2)
with equatorial semi axes a and b shown in Fig.1. For a spheroid with density $\rho$ angular velocity is given by:
$\Omega=\frac{2\pi}{P}=(2\pi\rho g(e))^{1/2},$ (3)
with
$g(e)=\frac{(1-e^{2})^{\frac{1}{2}}}{e^{3}}\arcsin(e)-\frac{3(1-e^{2})}{e^{2}}.$ (4)
The Maclaurin spheroid has angular momentum and mass defined as:
$J=I\Omega=\frac{2}{5}Ma^{2}\Omega,$ (5)
$M=\frac{4\pi}{3}\rho abc\approx\frac{4\pi}{3}\rho a^{3}(1-e^{2})^{1/2},$ (6)
where I is the moment of inertia, here due to the smallness of $\epsilon$ we take $a\approx b$. It was shown that for a equilibrium configuration with constant mass, entropy and chemical potential, the variation of energy, E, is related to the variation of angular momentum, J, according to Ostriker & Gunn (1969); Finn & Shapiro (1990) :
$dE=\Omega dJ.$ (7)
This relation can interpret the transformation of angular momentum into electromagnetic and gravitational energy emissions. During the spin down, eccentricity evolves from a maximum value to a small value, making the star less oblate. Maximum eccentricity for a Maclaurin spheroid, known as bifurcation point of Maclaurin and Jacobian sequence, is approximately 0.813 S.Chandrasekhar (1969), at this point bar-mode instabilities emerge. To avoid these dynamical instabilities, we consider maximum eccentricity value at bifurcation point.
In the following, we apply two different types of fluids i) incompressible and ii) compressible for the maclaurin spheroid to model magnetars as a central engine of GRBs.
Figure 1: A schematic view of an asymmetric rotating NS which rotates around $\hat{z}\rm-{axis}$, the ellipticity $\epsilon$ and eccentricity e are described by Eqs (1) and (2), respectively.
2.1 Incompressible model
Here, we introduce the incompressible Maclaurin spheroid and it’s evolutionary equations. In this model, the density of the fluid forming the spheroid is constant, therefore semi major axis, $a$, and moment of inertia, $I$, can be written as a function of eccentricity,
$a$ $=a_{0}(1-e^{2})^{-1/6},$ $a_{0}$ $=(\frac{3M}{4\pi\rho})^{1/3},$
$I$ $=I_{0}(1-e^{2})^{-1/3},$ $I_{0}$ $=\frac{2}{5}Ma_{0}^{2},$ (8)
where $a_{0}$ and $I_{0}$ are radius and inertial moment of a spherical object, respectively, assuming constant homogeneous density. According to these equations, the time evolution of the eccentricity induces a temporal changes in both $a$ and $I$. Assuming that mass remains unchanged during the evolution, one can find the following equation governing the variation of total angular momentum $J$ by taking the logarithmic derivative of Eqs (3), (5) and (2.1) :
$\frac{\dot{J}}{J}=\frac{1}{2}\left(\frac{g^{\prime}}{g}+\frac{4}{3}\frac{e}{1-e^{2}}\right)\dot{e},$ (9)
where $g^{{}^{\prime}}\equiv dg/de$. In next subsection, we discuss a more general form of maclaurin spheroid with polytropic density profile.
2.2 Compressible model
In order to consider compressible maclaurin spheroids, we take polytropic Equation of State (EoS) as
$P=K\rho^{\Gamma},$ (10)
where $P$ is pressure and $\Gamma=1+1/n$ and $n$ is the polytropic index. By solving Poisson’s equation for a uniformly rotating spheroid in hydrostatic equilibrium and applying proper symmetry and boundary conditions one can get Shapiro & Teukolsky (1983) :
$\left(\frac{\rho}{\rho_{0}}\right)^{\Gamma-4/3}=\frac{3}{2}(1-e^{2})^{2/3}A_{3}(e),$ (11)
Here $\rho_{0}$ is the constant density of a sphere ($e=0$) with radius $a_{0}$ and $A_{3}(e)$ is a function of the eccentricity which is given by:
$A_{3}(e)=\frac{2}{e^{2}}\left[1-(1-e^{2})^{1/2}\frac{\arcsin{e}}{e}\right].$ (12)
Note that in the limit of large $\Gamma$, compressible model of the fluid tends to incompressible model. Here, we have generalized form of set of equations (2.1) for compressible model:
$a$ $=a_{0}\left(\frac{\rho_{0}}{\rho}\right)^{1/3}(1-e^{2})^{-1/6},$
$I$ $=I_{0}\left(\frac{\rho_{0}}{\rho}\right)^{2/3}(1-e^{2})^{-1/3},$ (13)
Following the same strategy used for incompressible model and taking logarithmic derivative of Eq. (11) one can have:
$(\Gamma-2)\frac{d\rho}{\rho}=2\frac{da}{a}+\left(\frac{A^{\prime}_{3}}{A_{3}}-\frac{2e}{1-e^{2}}\right)de.$ (14)
Where $A_{3}^{{}^{\prime}}\equiv dA_{3}/de$. By substituting Eq. (14) into the logarithmic derivation of Eqs. (5) regarding $\dot{M}=0$, the following equation is obtained:
$\frac{\dot{J}}{J}=\frac{1}{2}\bigg{\{}\frac{g^{\prime}}{g}+\frac{1}{3\Gamma-4}\left[\frac{4(\Gamma-1)e}{1-e^{2}}-\frac{A^{\prime}_{3}}{A_{3}}\right]\bigg{\}}\dot{e}.$ (15)
Knowing the EM and GW emissions, through Eqs.(5) and (7) one can eliminate $J$ and $\dot{J}$ from Eqs. (9) and (15) and find the evolutionary equation for eccentricity for both compressible and incompressible models.
3 EM radiation and Spin down
In this section, we consider EM radiations from a magnetar to determine the evolution of eccentricity and then find the EM light curve. In subsection 3.1, we focus on dipole radiation, followed by calculations for quadrupole radiation in subsection 3.2, and the (dipole + quadrupole) total EM radiation in subsection 3.3.
3.1 Dipole luminosity
Energy loss via EM dipole moment (dipole luminosity) is given by:
$L_{\rm dip}$ $=-\dot{E}_{\rm dip}=\frac{2}{3}B_{\rm dip}^{2}a_{0}^{6}{\sin^{2}}{\alpha}\Omega^{4}$ (16)
$=8.13\times 10^{49}\rm erg/s\left(\frac{a_{0}}{10^{6}cm}\right)^{6}\left(\frac{0.5ms}{P}\right)^{4}\left(\frac{B}{10^{15}G}\right)^{2},$
where $P=2\pi/\Omega$, $B_{\rm dip}$ is the dipole magnetic field and $\alpha$ is the angle between the magnetic dipole moment and the rotational axis. hereafter, we select $\alpha=\pi/2$. In Eq. (16) we use appropriate values for scaling parameters.Combining Eqs. (16) and (7) the following equation for the angular momentum evolution is evaluated:
$\dot{J}_{dip}=-\frac{2}{3}B_{dip}^{2}a_{0}^{6}\Omega^{3}=-\beta\Omega^{3}.$ (17)
Finally putting Eqs. (5), (9) and (17) toghether, evolutionary equation for eccentricity in incompressible fluid is:
$\frac{de}{dt}=-\frac{\Omega^{4}(e)\beta}{\pi I_{0}\rho}(1-e^{2})^{1/3}\left[g^{\prime}(e)+\frac{4}{3}\frac{g(e)e}{1-e^{2}}\right]^{-1}.$ (18)
On the other hand, by using (5), (15) and (17) for the compressible fluid can obtain :
$\frac{de}{dt}$ $=-\frac{\Omega^{4}(e)\beta}{\pi I_{0}}\left(\frac{\rho_{0}}{\rho^{4}}\right)^{1/3}(1-e^{2})^{1/3}\times$ (19)
$\bigg{\{}g^{\prime}(e)+\frac{g(e)}{3\Gamma-4}\left[\frac{4(\Gamma-1)e}{1-e^{2}}-\frac{A^{\prime}_{3}}{A_{3}}\right]\bigg{\}}^{-1}.$ (20)
3.2 Quadrapole luminosity
Energy loss for EM quadrapole magnetic moment (quadrupole luminosity) is given by:
$\dot{E}_{\rm quad}=-\frac{32}{135}B_{\rm quad}^{2}a_{0}^{8}{\sin^{2}}{\beta_{1}}({\cos^{2}}{\beta_{2}}+10{\sin^{2}}{\beta_{2}})\Omega^{6},$ (21)
where $B_{\rm quad}$ is the quadrapole magnetic field and $\beta_{1}$ and $\beta_{2}$ are the inclination angles of the magnetic moments. It is instructive to introduce $\kappa$ as the quadrupole to dipole ratio:
$\kappa=\sin^{2}\beta_{1}\left(\cos^{2}\beta_{2}+10\sin^{2}\beta_{2}\right)\frac{B_{\rm quad}^{2}}{B_{\rm dip}^{2}}.$ (22)
An order of magnitude estimation for EM quadrapole luminosity is given by:
$L_{\rm quad}=\frac{32}{135}B_{\rm dip}^{2}a_{0}^{8}\kappa\Omega^{6}$ (23)
$=5.08\times 10^{51}erg/s\left(\frac{B}{10^{15}G}\right)^{2}\left(\frac{a_{0}}{10^{6}cm}\right)^{8}\left(\frac{0.5\rm ms}{P}\right)^{4}\left(\frac{\kappa}{500}\right).$
Therefore, angular momentum loss due to quadrapole magnetic moment radiations is:
$\dot{J}_{\rm quad}=-\xi\Omega^{5},$ (24)
where $\xi=-32/135B_{\rm quad}^{2}a_{0}^{8}\kappa$, leading to the evolutionary equation for the eccentricity of the incompressible model as:
$\frac{de}{dt}=-\frac{\Omega^{6}(e)\xi}{\pi I_{0}\rho}(1-e^{2})^{1/3}\left[g^{\prime}+\frac{4}{3}\frac{eg}{1-e^{2}}\right]^{-1}.$ (25)
For compressible model this evolutionary equation is evaluated as:
$\frac{de}{dt}$ $=-\frac{\Omega^{6}(e)\xi}{\pi I_{0}}(1-e^{2})^{1/3}\left(\frac{\rho_{0}}{\rho^{4}}\right)^{1/3}\times$ (26)
$\bigg{\{}g^{\prime}+\frac{g}{3\Gamma-4}\left[\frac{4(\Gamma-1)e}{1-e^{2}}-\frac{A^{\prime}_{3}}{A_{3}}\right]\bigg{\}}^{-1}.$
3.3 Total EM Luminosity
Now, we consider the total EM radiation in order to determine the angular momentum loss which is expressed as:
$\dot{J}_{\rm EM}=\dot{J}_{\rm dip}+\dot{J}_{\rm quad}=-\beta\Omega^{3}-\xi\Omega^{5}.$ (27)
Therefore, the following equations describe the evolution of eccentricity for the incompressible fluid as :
$\frac{de}{dt}=-\frac{\Omega^{4}(e)\beta}{\pi I_{0}\rho}(1-e^{2})^{1/3}\left(1+\frac{\Omega^{2}(e)\xi}{\beta\rho}\right)\left[g^{\prime}+\frac{4}{3}\frac{eg}{1-e^{2}}\right]^{-1},$ (28)
and in the case of the compressible model we obtain :
$\frac{de}{dt}$ $=\frac{-\Omega^{4}(e)\beta}{\pi I_{0}}(1-e^{2})^{1/3}\left(\frac{\rho_{0}}{\rho^{4}}\right)^{1/3}\left(1+\frac{\Omega^{2}(e)\xi}{\beta\rho}\right)\times$ (29)
$\bigg{\{}g^{\prime}+\frac{g}{3\Gamma-4}\left[\frac{4(\Gamma-1)e}{1-e^{2}}-\frac{A^{\prime}_{3}}{A_{3}}\right]\bigg{\}}^{-1}.$
Figure 2: Total EM luminosity for the incompressible model with fixed mass $M_{T}=2M\odot$. Different magnetic fields and quadrupole to dipole ratios are labeled.
Figure 3: EM dipole luminosity for the incompressible and compressible models, Here we assume that $M_{T}=2M\odot$ and $B=10^{15}G$. In the case of the compressible model, different adiabatic indexes are labeled.
Figure 4: EM dipole luminosity for the compressible model is plotted for $\Gamma=1.43$, where magnetic field and mass of the magnetar are labeled.
Figure 5: EM dipole and quadrupole emission of the compressible model with $\Gamma=1.43$, $B=10^{16}G$ and different fraction of dipole to quadrupole moments as labeled.
Fig. 2 shows the total EM luminosity for a magnetar containing incompressible fluid. Here, we assume that the dipole magnetic field strength is $B=10^{14}-10^{16}\rm G$ and the quadrupole is up to $500$ times larger than the dipole moment. It is seen that for higher ratios of magnetic quadrupole to dipole moments $\kappa$, the EM light curve of the magnetar starts to decay at earlier times with lower slope. This is a result of spin down with a faster rate at earlier times. We find that the luminosity is mainly powered by the dipole magnetic field at the earlier times and at later times it is dominated by dipole emission.
In Fig. 3, the effect of different polytropic equations of states (EoSs) on the dipole EM luminosity is investigated. It is illustrated that for some specific values of adiabatic index $\Gamma$, close enough to 4/3, a peak in the light curve appears. As it is seen, the peak fades away as $\Gamma$ increases and goes beyond $4/3$. These regions within the parameter space where the spin-up is caused by the decreasing moment of inertia due to the softness of the EoS, and black hole formation does not occur, we attribute the resulting flaring behavior in the light curve to these regions. The flaring behavior is closely tied to the rotational dynamics of new born NS which leads to a temporary spin-up of NS during its early spin-down phase. This effect crucially depends on the applied EoS and the adiabatic index $\Gamma$, as it is seen from Eq. (19) the non-axisymmetric deformations of the NS including a compressible fluid for $\Gamma\approx 4/3$ will show a non trivial behavior leading to a temporary spin-up behavior that are reflected in the observed flares upon the light curves.
The incompressible model is parameterized with $\Gamma=\infty$ and as it is seen from Fig. 3, it has the highest maximum initial luminosity for a given magnetic field and mass of the magnetar, and the maximum initial luminosity increases as softer EoSs are applied. Additionally, one can see that for stiffer EOSs, the decay time in the light curve occurs in the earlier times. Note that, the late time behavior of the light curve for all models with different $\Gamma$ values is similar. In Fig. 4, we explore the effect of both mass and magnetic field on the light curve of the dipole EM radiation for a compressible fluid with $\Gamma=1.43$. As shown in Fig. 3, the magnetic field has a significant impact on the peak luminosity, $L_{\rm p}$ and the peak time, $T_{\rm p}$, while the mass’s effect in a given range on $L_{\rm p}$ and $T_{\rm p}$ is not significant. The allowed range of the parameter space of mass and $\Gamma$ is determined through the stability analysis of the magnetar under deformation (see Appendix (A)). We introduce the valid mass range for our model based on black hole formation criteria following the method presented in Finn & Shapiro (1990); Wheeler & Klauder (1972)
In Fig. 5, the magnetic dipole luminosity is compared to the total EM luminosity including quadrupole emission with different $\kappa$ in the compressible model. We find that taking the quadrupole magnetic moment into account lead to an earlier peak time in the light curves. Moreover, for a given dipole magnetic field, the effect of the quadrupole to dipole moment ratio on the peak luminosity is negligible.
4 GW radiation and Spin down
4.1 GW luminosity
In this section, we investigate the GW radiations from a deformed magnetar due to the asymmetry in its equatorial plane. The GW luminosity of a deformed magnetar is expressed as Andersson (2019):
$L_{\rm GW}=-\dot{E}_{\rm GW}=\frac{1}{5}\left<\frac{d^{3}{I}_{jk}}{dt^{3}}\frac{d^{3}{I}_{jk}}{dt^{3}}\right>=\frac{32}{5}I^{2}\epsilon^{2}\Omega^{6}$
$=1.80\times 10^{53}\,\text{erg/s}\left(\frac{a_{0}}{10^{6}\,\text{cm}}\right)^{4}\left(\frac{\epsilon}{10^{-3}}\right)^{2}\left(\frac{0.5\,\text{ms}}{P}\right)^{6}\left(\frac{M}{2.6M_{\odot}}\right)^{2}.$ (30)
By introducing $\gamma_{0}\equiv 32I_{0}^{2}\epsilon^{2}/5$, the time evolution of the eccentricity due to GW radiation from an incompressible fluid is given by the following equation :
$\frac{de}{dt}=-\frac{\Omega^{6}(e)\gamma_{0}}{\pi I_{0}\rho}(1-e^{2})^{-1/3}\left[g^{\prime}(e)+\frac{4}{3}\frac{g(e)e}{1-e^{2}}\right]^{-1}.$ (31)
and for the compressible fluid one has :
$\frac{de}{dt}$ $=-\frac{\Omega^{6}(e)\gamma_{0}}{\pi I_{0}\rho}\left(\frac{\rho_{0}}{\rho}\right)^{5/3}(1-e^{2})^{-1/3}\times$ (32)
$\bigg{\{}g^{\prime}(e)+\frac{g(e)}{3\Gamma-4}\left[\frac{4(\Gamma-1)e}{1-e^{2}}-\frac{A^{\prime}_{3}}{A_{3}}\right]\bigg{\}}^{-1}.$
Figure 6: GW luminosity for compressible model for $\Gamma=1.43$, and $\epsilon=10^{-3}$ for different masses of NS as labeled.
Figure 7: GW luminosity for compressible model for $\Gamma=1.53$, and $\epsilon=10^{-3}$ for different masses of NS as labeled.
Figure 8: GW luminosity for incompressible ($\Gamma=\infty$) and compressible models, Here we assume that $M_{T}=2M\odot$ and $\epsilon=10^{-3}$, in the case of compressible model different adiabatic indexes ($\Gamma$) are labeled.
We plot $L_{GW}$ for some cases in Fig. 6 - 8 to examine the gravitational energy loss from a newborn magnetar. Fig. 6 and Fig. 7 describe GW luminosity for the compressible model for $\Gamma=1.43$ and $\Gamma=1.53$ respectively.These light curves indicate that for higher masses, the luminosity peak appears at earlier times with higher magnitude. At late times light curves for all masses coincide. Comparing Fig. 6 and Fig. 7, light curve peaks in Fig. 6 are more apparent. As a result, a higher frequency signal should be detected for a magnetar with softer EOS during peak time. Fig. ‘8 demonstrates GW luminosity for different EOSs. As $\Gamma$ tends to $4/3\approx 1.33$, the peak in the light curves fades away, which is the same result as seen in the EM light curve. Furthermore, the behavior of maximum luminosity with respect to changes in $\Gamma$ is the same.
4.2 Detectability of the GW Signals
Here, we consider the detectiblity of the GW signal originating from a deformed magnetar in the light of future GW detectors. The GW strain for the deformed magnetar located at distance d from the observer is given by Howell et al. (2011); Bambi et al. (2021); Corsi & Mészáros (2009):
$h(t)=$ $\frac{4}{d}{\epsilon}I\Omega(t)^{2}\ \approx 5.31\times 10^{-25}\times$ (33)
$\left(\frac{40Mpc}{d}\right)\left(\frac{a_{0}}{10^{6}cm}\right)^{2}\left(\frac{\epsilon}{10^{-3}}\right)\left(\frac{0.5ms}{p}\right)^{2}\left(\frac{M}{2.6M_{\odot}}\right).$
The characteristic GW strain as a dimensionless quantity can be calculated using
$h_{c}=fh(t)\sqrt{dt/df}.$ (34)
Knowing that $f=\Omega/\pi$ one can find that:
$h_{c}=\frac{4\pi^{2}}{d}{\epsilon}f^{3}\sqrt{dt/df}.$ (35)
In order to find the explicit form of $h_{c}$ one needs to find the term $dt/df$, which can be find according to:
$\frac{dE_{GW}}{df}=\frac{dE_{GW}}{dt}\frac{dt}{df}=L_{GW}\frac{dt}{df}.$ (36)
Considering the domination of GW radiation at early times:
$\frac{dE}{df}\approx\frac{dE_{GW}}{df}=\pi^{2}If,$ (37)
which leads to the following characteristic GW strain as:
$h_{c}(t)=\frac{1}{d}\sqrt{\frac{5If}{2}}=\frac{1}{d}\sqrt{\frac{5I\Omega(t)}{2\pi}}$ (38)
$\approx 1.45\times 10^{-21}\left(\frac{40Mpc}{d}\right)\left(\frac{a_{0}}{10^{6}cm}\right)\left(\frac{0.5ms}{P}\right)^{1/2}\left(\frac{M}{2.6M_{\odot}}\right)^{1/2}.$
In order to find $h(t)$ one needs to determine $\Omega(t)$ due to variation of eccentricity, we plot $h_{c}$ in Fig. 9 for different masses as labled assuming $d=40Mpc$, $a_{0}=10^{6}cm$ and $P=0.5$ ms for a new born magnetar. It is seen than the variation of the mass in the given range has not significant impact on $h_{c}$ and the resulting curves for three different values of masses almost coincide each other.
On the other hand, for later times, EM radiation dominates the energy loss, by rewriting Eqs. (36) and (37) for EM case, $dt/df$ is given as bellow:
$\frac{dt}{df}\approx\frac{\pi^{2}If}{L_{EM}},$ (39)
the characteristic GW strain for this regime is obtained as:
$h_{c}(t)$ $=\frac{1}{d}\left(1+\frac{16}{45}\pi^{2}a^{2}\kappa\Omega(t)^{2}\right)^{-1/2}\sqrt{\frac{24I^{3}\epsilon^{2}\Omega(t)^{3}}{\pi B_{dip}^{2}a^{6}}},$ (40)
Here, we can ignore the effect of quadruple magnetic field while $\kappa a^{2}\Omega^{2}<<1$, and write $h_{c}(t)$ only for EM dipole radiations:
$h_{c}(t)=$ $\frac{1}{d}\sqrt{\frac{24I^{3}\epsilon^{2}\Omega(t)^{3}}{\pi B_{dip}^{2}a^{6}}}\approx 9.54\times 10^{-22}$ (41)
$\left(\frac{40Mpc}{d}\right)\left(\frac{0.5ms}{P}\right)^{3/2}\left(\frac{M}{2M_{\odot}}\right)^{3/2}\left(\frac{\epsilon}{10^{-3}}\right)\left(\frac{10^{16}G}{B_{\text{dip}}}\right).$
Characteristic GW strain for the case in which the spin down drives by the EM dipole radiation is plotted in Fig. 9 for various $B$ and $\epsilon$ as labeled. We see from Eq. (41) and Fig. 9 that a large ellipticity would produce a stronger gravitational wave signal. Eq. (41) is inversely proportional to the $B_{\text{dip}}$, however the magnetic field should be large enough to reproduce the observed EM luminosity therefore magnetic field strength has a lower bound. On the other hand it was shown that the large ellipticity can be generated thanks to high values of the magnetic field, which means larger magnetic field components Xie et al. (2022b); Sur & Haskell (2021); Lasky & Glampedakis (2016); Stella et al. (2005).
The projected sensitivities for aLIGO O5, Einstein Telescope (ET) and Cosmic Explorer (CE) detectors are also presented in Fig. 9, it is found that GW signals from magnetars with the specific set of parameters, reach the sensitivity threshold of theses detectors. However, being $h_{c}$ within the sensitivity range of GW detector (as it is seen from Fig. 9) does not necessarily guarantee that the GW will be detected, though it suggests that detection should be possible under ideal conditions. In practice, the detector’s ability to identify a GW signal depends on several factors, such as noise, duration of the signal, and the methods used to search for the signal. If the signal to noise ratio (SNR) is too low (e.g., the signal is weak or buried in background noise), it may not be detectable. For continuous or long-duration signals (such as those expected from deformed neutron stars or magnetars), noise can accumulate over time, and the signal can be buried in background noise, making detection harder despite being within the sensitivity range. Imperfect search algorithms or unknown signal characteristics could also result in a missed detection. Therefore, being in the sensitivity range is a necessary but not sufficient condition for the detection. However, if the GW signal associated with the EM light curve of the magnetar can be detected in the future, it would be a strong evidence that the magnetar can act as the central engine of GRBs.
Figure 9: Sensitivities of GW detectors to GW characteristic strain produced by rotating magnetars for different masses as labeled.
[!htbp]
5 GW and EM spin down
This section describes the entire energy loss of the magnetar during EM and GW emissions. Despite the fact that the total luminosity is not an observable quantity by a single instruments, it can illustrate the dynamical behavior of a new born magnetar through multi-messenger emissions.
5.1 Dipole EM and GW Radiations
Here, we consider spin down due to both dipole and GW radiations. The resulting evolutionary equation for eccentricity in incompressible is given by:
$\frac{de}{dt}=$ (42)
$-\frac{\Omega^{4}(e)\beta}{\pi I_{0}\rho}(1-e^{2})^{1/3}\left[1+\frac{\Omega^{2}(e)\gamma_{0}}{\beta(1-e^{2})^{2/3}}\right]\left[g^{\prime}(e)+\frac{4}{3}\frac{g(e)e}{1-e^{2}}\right]^{-1},$
and for the compressible model, we have;
$\frac{de}{dt}=$
$-\frac{\Omega^{4}(e)\beta}{\pi I_{0}}\left(\frac{\rho_{0}}{\rho^{4}}\right)^{1/3}(1-e^{2})^{1/3}\left[1+\left(\frac{\rho_{0}}{\rho}\right)^{4/3}\frac{\Omega^{2}(e)\gamma_{0}}{\beta(1-e^{2})^{2/3}}\right]\times$
$\bigg{\{}g^{\prime}(e)+\frac{g(e)}{3\Gamma-4}\left[\frac{4(\Gamma-1)e}{1-e^{2}}-\frac{A^{\prime}_{3}}{A_{3}}\right]\bigg{\}}^{-1}.$ (43)
Figure 10: Comparing dipole,GW and dipole+GW luminosity for compressible model, Here we assume that $M_{T}=2M\odot$, $B=5\times 10^{15}G$ and $\epsilon=10^{-3}$.
Fig. 10 shows the resulting luminosity from EM dipole and GW radiations for incompressible model of a magnetar with $M_{T}=2M\odot$ and assuming $B=5\times 10^{15}G$ and $\epsilon=10^{-3}$, the total luminosity (EM+GW) is also represented in this figure. According to this plot, it is seen that EM luminosity has a sharper peak which appears with a delay after GW peak time and at the early times GW luminosity is dominated. For a time interval around dipole EM peak, it dominates and then decays with a faster rate than GW luminosity. The total luminosity on the other hand starts to decay earlier and faster with a slope larger than dipole luminosity’s and less than GW one.
It is instructive to compare our results with those presented in Corsi & Mészáros (2009), where it has been shown that the peak amplitude of the GW signal occurs with a delay compare to the GRB trigger, in our scenario we see that for a given set of parameters, the GW flares can also even appear prior to EM ones.
5.2 Dipole and Quadrapole EM and GW Radiations
Taking the quadrapole magnetic moment radiation into our consideration, the evolutionary equations for eccentricity in incompressible and compressible models are given by:
$\frac{de}{dt}$ $=\frac{-\Omega^{4}(e)\beta}{\pi I_{0}\rho}(1-e^{2})^{1/3}\times$ (44)
$\bigg{\{}1+\left[\xi+\frac{\gamma_{0}}{(1-e^{2})^{2/3}}\right]\frac{\Omega^{2}(e)}{\beta}\bigg{\}}\left[g^{\prime}+\frac{4}{3}\frac{eg}{1-e^{2}}\right]^{-1},$
and
$\frac{de}{dt}$ $=$ (45)
$\frac{-\Omega^{4}(e)\beta}{\pi I_{0}}(1-e^{2})^{1/3}\left(\frac{\rho_{0}}{\rho^{4}}\right)^{1/3}\times$
$\left[1+\bigg{\{}\xi+\gamma_{0}\left(\frac{\rho_{0}}{\rho}\right)^{4/3}(1-e^{2})^{-2/3}\bigg{\}}\frac{\Omega^{2}(e)}{\beta}\right]\times$
$\bigg{\{}g^{\prime}+\frac{g}{3\Gamma-4}\left[\frac{4(\Gamma-1)e}{1-e^{2}}-\frac{A^{\prime}_{3}}{A_{3}}\right]\bigg{\}}^{-1}.$
Figure 11: GW, EM (dipo+quad) and GW+EM luminosities for the incompressible model assuming $\epsilon=10^{-3}$ and $\kappa=100$ for total mass $M=1.6M\odot$ and $M=2M\odot$ as labeled.
Figure 12: Comparing EM (dipo+quad), GW and total luminosity for a $2M\odot$ magnetar modeled with compressible fluid applying $\Gamma=1.43$ and $\epsilon=10^{-3}$, (top) assuming $B=10^{15}G$ and $\kappa=50$, (bottom) $B=5\times 10^{15}G$, $\kappa=100$.
Fig. 11 illustrates a comparison of the EM (dipo+quad), GW and the total (GW+EM) luminosity for the incompressible model of nascent NSs with $M=1.6M\odot$ and $M=2M\odot$ regarding $\epsilon=10^{-3}$, $\kappa=100$ and $\Gamma=1.43$. We see that the GW luminosity dominates at earlier times and EM radiation overcomes at later times. Similar to the behavior observed in Fig. 10, the total radiation luminosity decays with a faster rate. In Fig. 12, the behavior of EM, GW and total luminosity are illustrated for two different set of parameters. In both the top and bottom panels of Fig. 12, a compressible magnetar with $\Gamma=1.43$ and $\epsilon=10^{-3}$ is assumed, but $B$ and $\kappa$ is different. This figure shows the crucial dependence of the appearance time of EM or GW peaks, in the top panel for $B=10^{15}G$ and $\kappa=50$, the GW peak luminosity occurs perior to the EM peak luminosity, while in the bottom panel the priority of these peaks are changed by assuming $B=5\times 10^{15}G$, $\kappa=100$. Our investigation in this section shows that in the case of appearence of peak within the luminosity there is a delayed arrival time between peak in GW and EM radiations which provide a unique signature can be confirmed by ongoing and upcoming multimessenger observations of high energy astrophysics satellites and ground based detectors.
6 FITting GRB Flares within compressible magnetar model
The GRB afterglow observed in different energy bands from X-ray to the optical and radio bands which follows a similar power-law, but with significant higher luminosity in X-ray band. Consequently, the X-ray luminosity can be served as an indicator of the total (bolometric) afterglow luminosity. It is worth mentioning that due to the anisotropic radiation effects of GRBs introduced by the jet structure and also the instrumental limitations, the bolometric luminosity might be larger than the observed luminosity. Since Swift-XRT generally offers the most comprehensive observations of afterglow evolution, in this paper we rely on the afterglow luminosity data captured by Swift-XRT in order to compare with our model predictions. In some of the previous studies, the bolometric luminosity is estimated by multiplying a factor of 2 - 7 to the energy contained in the Swift-XRT band (0.3-10 keV) Rowlinson et al. (2013); Rueda et al. (2022); Wang et al. (2024).
Here, we have focused on a GRB sample including GRBs 070306, GRB 090812, GRB 100513A and GRB 100901A, which all sources have known redshifts, and their modified light curves presented using k-correction (Bloom et al., 2001). All the GRB sources in our sample have the maximum luminosity of the X-ray flares less than $10^{51}\text{erg s}^{-1}$ and the peak of the luminosity occur in less than 200 s. In order to model the X-ray flares, we consider a new born NS with $2M_{\odot}$ and free parameters are assumed to be the magnetic field $B$, the quadrupole to dipole ratios $\kappa$, and adiabatic index $\Gamma$. Here, we adopt Metropolis-Hasting (M-H) algorithm from Markov Chain Monte Carlo (MCMC) method in Figs. 14 - 16 to find the best fit for deformed newborn magnetar model (presented in sections 2 and 3) taking into account the Swift XRT data. The MCMC python code is also employed in order to generate posterior samples, we plotted a 2d distribution for posteriors in the right pannels of Figs. 14 - 16 showing pairwise correlation between parameters. In order to find the best fit parameters, we used Swift/XRT data for X-flares occurring in less than few hundred seconds (colored purple data points in Figs. 14 - 16) which covers the peak feature of X-ray flares. For each parameters ($B$, $\kappa$ and $\Gamma$), we estimated probability distribution function (PDF) with gaussian kernel density estimation (KDE) on sample’s histogram given in the right pannels of Figs. 14 - 16. The value for width of gaussian function is chosen to be 0.5, showing that the range of valid parameters to fit the data. The dashed magenta lines in PDF plots indicate the median of each parameter sample. We provide $\chi^{2}_{\text{red}}=\chi^{2}/dof$ (chi-squared divided by the degrees of freedom) in Table. 1, and all the fitting parameters summarized in this table with the errors are given as well. The obtained values of fitting parameters shows that $B\sim(6-20)\times 10^{15}G$, $\kappa\sim 49-150$ and $\Gamma\sim 1.5-1.7$ in order to explain the X-ray flares occur in less than few handered seconds after the Swift trigger time. The model presented in this paper is limited to the Maclaurian spheroids, however this investigation can be extended to the Jacobi ellipsoid, in the case that the eccenterity is larger than 0.81 and also for the case in which 3 axial are different.
We find that the peaks in Swift data can be produced by a softer EOS in which $\Gamma$ tends to $1.43$, note that the plateau behaviour in later times is given by stiffer EOSs. Generally magnetic field, $\kappa$ and EOS are not constant thorough dynamical evolution of deformed magnetars, these features should be taken into account in a more realistic configurations, and will be the subject of a future paper. It is worth mentioning to note that while in the present paper, we focus on the modeling of GRB X-ray flares by compressible new born magnetars, our model has the potential to explain the X-ray plateau (in addition to X-ray flares) taking into account different types of the fluids with multipolar structures of the magnetic field Wang et al. (2024); Xie et al. (2022b). We plan to perform a comprehensive statistical analysis for a large sample of X-ray flares in a future paper. Taking into account the time dependency of the magnetic field and also including accretion materials onto a new born magnetar could be origin of the X-ray flares in GRB light curves and can be considered a more generalized model.
7 Conclusions
In this paper, we presented a scenario based on dynamical evolution of a Maclaurin spheroid to model GRB’s central engine taking into account both compressible and incompressible fluids in order to explain different features of EM and GW luminosities. We evaluated the effect of the total mass $M_{T}$, the magnetic field $B$, the ratio of the magnetic quadrupole to dipole moment $\kappa$ and the adiabatic index $\Gamma$ on the light curves of a deformed NS. We showed that the general properties of the EM/GW light curve crucially depends on the nature of EOS which is applied to model the magnetar and we found that for a given set of parameters, our model can generate flaring activities in both GW and EM luminosities. Our main motivation inspired by Finn & Shapiro (1990), where the rotational evolution of a new NS is originating from the deformation of the object which can result in both spin down and spin up motion in various circumstances. Besides considering the EM signal of our GRB model, the GW signal and its detectability in the light of the sensitivity range of GW detectors has been considered. In Sec. 6 for a sample of GRBs by using Swift/XRT data, we applied MCMC method in order to find the best fit parameters of our model. In this study, we tried to provide valuable insights into the underlying mechanisms driving the observed temporal behaviors in GRB afterglows such as the presence of X-ray flares based on a deformed magnetar model, in this regard various multi-messenger predictions of the model have been presented and also the validity of the model has been examined in the light of some of the Swift/XRT observational data.
Figure 13: In the left panel, the compressible model for a magnetar with $M=2M_{\odot}$ is used to fit the Swift/XRT data of GRB 100901A. The colored purple data points (which covers the X-ray flares) are applied to find the best fit model parameters $B_{\text{dip}}$, $\kappa$ and $\Gamma$. In the right panel, the posterior distributions for fitting parameters are plotted taking into account the histogram and estimated PDF for each parameter. The dashed magenta lines in PDF plots illustrate the median of the parameters, and the best fit values are represented at the top the diagrams.
Figure 14: The same as Fig. 13, but for GRB 090812.
Figure 15: The same as Fig. 13, but for GRB 100513A.
Figure 16: The same as Fig. 13 but for GRB 070306.
GRB Name Redshift $L_{\text{peak}}(\mathrm{erg\ s^{-1}})$ $t_{\text{peak}}(\mathrm{s})$ B (G) $\Gamma$ $\kappa$ $\chi_{\text{red}}^{2}$
GRB 070306 1.496 $1.03\times 10^{50}$ 73.56 $1.12^{+0.03}_{-0.03}\times 10^{16}$ $1.59^{+0.00}_{-0.00}$ $142.8^{+1.58}_{-1.58}$ 2.43
GRB 090812 2.452 $2.61\times 10^{50}$ 58.14 $2.01^{+0.04}_{-0.04}\times 10^{16}$ $1.57^{+0.03}_{-0.03}$ $49.6^{+0.65}_{-0.65}$ 1.16
GRB 100513A 4.772 $7.16\times 10^{49}$ 35.49 $7.23^{+0.84}_{-0.84}\times 10^{15}$ $1.70^{+0.07}_{-0.07}$ $116.1^{+4.45}_{-4.45}$ 0.93
GRB 100901A 1.408 $2.25\times 10^{49}$ 175.59 $6.89^{+0.79}_{-0.79}\times 10^{15}$ $1.50^{+0.01}_{-0.01}$ $148.0^{+3.33}_{-3.33}$ 1.27
Table 1: Constraints on properties of the deformed newborn magnetar via fitting the X-ray light curve of GRBs. The table presents the fitting results, including the magnetic field strength (B), the adiabatic index $\Gamma$, and the ratio of the magnetic quadrupole to dipole moment ($\kappa$).
Acknowledgments
S.S and P.H would appreciate for valuable discussions during the conferences of High Energy Astrophysics and Cosmology in the era of all-sky in October 2024, Yerevan, Armenia. S.S and P.H are grateful to Massimo Della Valle for insightful comments and discussions.
Appendix A
Stability analysis and Black Hole formation
In this appendix, we consider the criteria of the black hole (BH) formation for both incompressible and compressible models in order to determine allowed region of the parameter space over the mass and adiabatic index. For the incompressible model, we should evaluate the Schwarzschild radius at $a_{0}=a(e=0)$ to investigate whether the BH is formed or not. But in the case of compressible model, semi major axis $a$ is a more complicated function of the eccentricity and it is not obvious that if $a_{0}$ gives the minimum radius. Taking into account the logarithm of Eq. (11), we find that :
$(\Gamma_{\text{crit}}-\frac{4}{3})\log\left(\frac{\rho_{\text{crit}}}{\rho_{0}}\right)=\frac{3}{2}(1-e^{2})^{2/3}A_{3}(e).$ (A1)
By considering the critical value for the density as below
$\rho_{\text{crit}}=\frac{M}{(4\pi/3)a_{\text{crit}}^{3}}=\frac{(1-e^{2})^{1/2}}{(32\pi/3)M^{2}},$ (A2)
where $a_{\text{crit}}\sim R_{S}=2M$. Therefore, one can find a critical value for $\Gamma$ as (see Eq. 63 Finn & Shapiro (1990))
$\Gamma_{\text{crit}}=\frac{\log[2(1-e^{2})^{-2/3}/3A_{3}(e)]}{\log[(32\pi/3)M^{2}\rho_{0}(1-e^{2})^{-1/2}]}+\frac{4}{3}.$ (A3)
For a NS with mass M if $\Gamma<\Gamma_{\text{crit}}$ at a given value of eccentricity, the star will collapse to a BH eventually. The stability analysis of the NS are displayed in Figs. 17 and 18. In left panel of Fig. 17, we present $\Gamma_{\text{crit}}$ for various masses in range $(1.6-2.6)M_{\odot}$ as a function of eccentricity. For all chosen masses, all $\Gamma_{\text{crit}}$ curves lies below $\Gamma=1.43$ line at the given range of eccentricity meaning that a NS with a mass within this range remains stable during its evolution. On the right panel of Fig. 17, the mass range $(3.6-3.9)M_{\odot}$ is adopted, within this range all $\Gamma_{\text{crit}}$ curves are above $\Gamma=1.43$ line at the beginning of evolution (for larger e) and they cross the horizontal line at some eccentricity values. This means that for a range of M and e, we can not have stable configuration for $\Gamma=1.43$ and therefore this part of parameter space is excluded in our consideration.
In Fig. 18, we present a density plot for $\Gamma_{\text{crit}}$ as a function of mass and eccentricity. The presence of a singularity in the denominator of Eq. (A3) for some values of mass and eccentricity leads to infinite value for $\Gamma_{\text{crit}}$, showing in the plot with the sharper color. Regarding the discussion mentioned above, this plot provide essential information about the maximum allowed value of $\Gamma$ and therefore the stable region of our parameter space.
Figure 17: The plot shows the variation of the adiabatic index for various masses as labeled, The colored lines show $\Gamma_{\text{crit}}$ for different masses and the horizontal dashed lines display $\Gamma=1.43$.
Figure 18: The density plot $\Gamma_{\text{crit}}$ is displayed as a function of both mass and eccentricity. This figure contains essential information about the stability of NS which is plotted in log scale.